# The extension of Schwarzschild line element to include uniformly accelerated mass

### v3

@@ B.3 Rindler transformation for the frames with non-zero initial velocity
$\text{or\,\,\,\,\,\,}\bar{x}$ $=\Bigg{[}\left(x+\gamma\frac{c^{2}}{a_{x}}\right)^{2}-\left(ct+\gamma\frac{v_{x}c}{a_{x}}\right)^{2}\Bigg{]}^{1/2}-\frac{c^{2}}{a_{x}}$ (B.37)
Therefore the inverse Rindler transformation for an accelerated frame having non-zero initial velocity can be written by using equations (B.36) and (B.37) as follows:
$c\bar{t}$ $=\frac{c^{2}}{a_{x}}\tanh^{-1}\left\{\frac{\left(ct+\gamma\frac{v_{x}c}{a_{x}}\right)-\frac{v_{x}}{c}\left(x+\gamma\frac{c^{2}}{a_{x}}\right)}{\left(x+\gamma\frac{c^{2}}{a_{x}}\right)-\frac{v_{x}}{c}\left(ct+\gamma\frac{v_{x}c}{a_{x}}\right)}\right\}$ (B.38)
$\bar{x}$ $=\Bigg{[}\left(x+\gamma\frac{c^{2}}{a_{x}}\right)^{2}-\left(ct+\gamma\frac{v_{x}c}{a_{x}}\right)^{2}\Bigg{]}^{1/2}-\frac{c^{2}}{a_{x}}$ (B.39)
$\bar{y}$ $=y$ (B.40)
$\bar{z}$ $=z$ (B.41)

### v4

@@ A.3 Rindler transformation for the frames with non-zero initial velocity
$\text{or\,\,\,\,\,\,}\bar{x}$ $=\Bigg[\left(x+\gamma\frac{c^{2}}{a_{x}}\right)^{2}-\left(ct+\gamma\frac{v_{x}c}{a_{x}}\right)^{2}\Bigg]^{1/2}-\frac{c^{2}}{a_{x}}$ (A.37)
Therefore the inverse Rindler transformation for an accelerated frame having non-zero initial velocity can be written by using equations (A.36) and (A.37) as follows:
$c\bar{t}$ $=\frac{c^{2}}{a_{x}}\tanh^{-1}\left\{\frac{\left(ct+\gamma\frac{v_{x}c}{a_{x}}\right)-\frac{v_{x}}{c}\left(x+\gamma\frac{c^{2}}{a_{x}}\right)}{\left(x+\gamma\frac{c^{2}}{a_{x}}\right)-\frac{v_{x}}{c}\left(ct+\gamma\frac{v_{x}c}{a_{x}}\right)}\right\}$ (A.38)
$\bar{x}$ $=\Bigg[\left(x+\gamma\frac{c^{2}}{a_{x}}\right)^{2}-\left(ct+\gamma\frac{v_{x}c}{a_{x}}\right)^{2}\Bigg]^{1/2}-\frac{c^{2}}{a_{x}}$ (A.39)
$\bar{y}$ $=y$ (A.40)
$\bar{z}$ $=z$ (A.41)